OpenStudy (mckenzieandjesus):

Find the common ratio of the sequence. –164, –82, –41, –20.5, . . .

jimthompson5910 (jim_thompson5910):

Pick any term you want except for the first term. Divide this term by the previous term.

OpenStudy (mckenzieandjesus):



OpenStudy (mckenzieandjesus):

oh okay i did it a different way and couldn't figure out what 10 1/4th would be so i just clicked 1/2.. comes to find out 2-20.5/-41= 1/2 :)

jimthompson5910 (jim_thompson5910):

what that work shows is that you're dividing each term by 2 to get the next term notice how (2nd term)/(1st term) = (-82)/(-164) = 1/2 = 0.5

jimthompson5910 (jim_thompson5910):

that's an interesting way to get the answer

OpenStudy (mckenzieandjesus):

Yeah thats what I figured out above :)

jimthompson5910 (jim_thompson5910):

`so 10 1/4 does = 1/2?` idk what you mean here

jimthompson5910 (jim_thompson5910):

10 1/4 is not equal to 1/2

jimthompson5910 (jim_thompson5910):

but yeah the common ratio is 1/2 or 0.5

jimthompson5910 (jim_thompson5910):



OpenStudy (mckenzieandjesus):

ohh

OpenStudy (mckenzieandjesus):

well I still clicked the right answer :)

OpenStudy (mckenzieandjesus):

thank u for showing me the correct way to do it